Circular No. 7441

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cp IN PGC 57064 M. Modjaz, S. Beckmann, C. Yu, and W. D. Li, University of Califor-

nia at Berkeley, on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope of an apparent supernova in the Abell cluster 2151. SN 2000cp (mag ∼ 18.0) was discovered on an unfiltered image taken on June 21.3 UT, and was confirmed on an earlier image taken on June 15.3 (mag ∼ 18.8). SN 2000cp is located at α = 16^{h}05^{m}27^{s}.62, δ = +17^{o}49′48′′.8 (equinox 2000.0), which is between the two apparent nuclei of PGC 57064 and is 3′′.3 west and 1′′.7 north of the southern nucleus. An unfiltered image of the field taken on June 11.3 showed nothing at the position of SN 2000cp (limiting mag ∼ 19.0).

AX J0057.4–7325 AND SMC X-2 K. Torii, National Space Development Agency of Japan; T. Kohmura,

Osaka University; and J. Yokogawa and K. Koyama, Kyoto University, communicate: “An ASCA observtion made on Apr. 25–26 detected coherent pulsations of period 101.42 ± 0.06 s from a new source in the Small Magellanic Cloud, designated AX J0057.4–7325 and located at α = 0^{h}57^{m}27^{s}.0, δ = −73^{o}25′31′′ (equinox 2000.0; error radius 1′). The spectrum is characterized by a flat power-law function with photon index 0.9 and x-ray flux 2.4 × 10^{−12} erg cm^{−2} s^{−1} (0.7–10 keV). Within the error region is located a ROSAT source, RX J0057.3–7325 (Kahabka et al. 1999, A.Ap. Suppl. 136, 81). Coherent pulsations of period 2.37230 ± 0.00004 s from SMC X-2, a 2.37-s pulsar (IAUC 7402), were also found in the same observation. The photon index and x-ray flux of SMC X-2 were 0.6 and 1.0 × 10^{−11} erg cm^{−2} s^{−1} (0.7–10 keV), respectively.”

SUPERNOVA 2000cn IN UGC 11064 K. Krisciunas, University of Washington; and F. Deglman and R.

McMillan, Apache Point Observatory (APO), report preliminary photometry (accurate to ± 0.02 mag) of SN 2000cn (IAUC 7436), using the APO 3.5-m telescope: June 12.24 UT, V = 16.63, B − V = +0.18, V − R = +0.15, V − I = +0.05; June 16.23, 16.68, +0.35, +0.13, −0.07.

2000 June 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green